40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire Unknown.jpeg
    Reply  Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire 
    Reply  Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp e12342.jpg
    Reply  Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test 43micr242grs.jpg
       Reply  Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck 
          Reply  Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck 
             Reply  Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp 
                Reply  Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw 
          Reply  Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended 
             Reply  Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
             Reply  Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended 
                Reply  Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
                   Reply  Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 
                   Reply  Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
                      Reply  Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful ETMX_resonances.pdfETMX_SUSspectra.zip
                         Reply  Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up 
                            Reply  Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up IMG_2996.JPGIMG_2997.JPG
                               Reply  Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up 
                                  Reply  Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
                                     Reply  Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended 
                                        Reply  Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended pitchBalancingWithBias.PDF
                                           Reply  Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation IMG_6816.JPG
                                              Reply  Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge AlBridge.jpg
                                              Reply  Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS IMG_3005.JPG
                                                 Reply  Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished 
                                                    Reply  Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers IMG_3006.JPG
                                                       Reply  Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation 800P.jpg
                                                       Reply  Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished 
                                                          Reply  Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted 
                                                             Reply  Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended IMG_3018.JPG
                                                                Reply  Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended IMG_3019.JPG
                                                                   Reply  Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted 
                                                                      Reply  Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
                                                                         Reply  Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO ETMY_BounceSpectra_25Aug2016.pdf
                                                                            Reply  Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_BounceSpectra_26Aug2016.pdf41.png19.png
                                                                               Reply  Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic  ETMY-ULhistory.png
                                                                               Reply  Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic  
                                                                                  Reply  Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_UL_problematic.png
                                                                                     Reply  Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic  satelliteBox.png
                                                                                        Reply  Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics IMG_6840.JPGIMG_6841.JPG
                                                                            Reply  Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction ETMY_BounceSpectra_26Aug2016_1.pdf
                                                                               Reply  Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling IMG_3051.JPG42.png
                                                                                  Reply  Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed IMG_3052.JPGSRM_sensor_level_comp.pngETMY_BounceSpectra_30Aug2016.pdf
                                                                                     Reply  Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging 
                                                                                        Reply  Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green  
                                                                                           Reply  Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick ITMY_SENSOR_UL.pngspareSatAmp.jpg
                                                                                              Reply  Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick 
                                                                                                 Reply  Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned 
                                                                                           Reply  Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm 
                                                                                              Reply  Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 03.png56.png
                                                                                                 Reply  Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber IMG_3137.JPGIMG_3143.JPGIMG_3142.JPGIMG_3148.JPG
                                                                                                    Reply  Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible 
                                                                                                       Reply  Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers 
                                                                                                          Reply  Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment Osamu.JPG
                                                                                                             Reply  Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment IMG_3273.JPGIMG_3275.JPG
                                                                                        Reply  Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
                                                                                           Reply  Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling 
                                                                                     Reply  Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments 
                                                                                        Reply  Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments 
Message ID: 12377     Entry time: Fri Aug 5 02:17:10 2016     In reply to: 12370     Reply to this: 12378   12381
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMX Ruby Wire Standoff Ready for Gluing 

Part 1: Rotation of optic

  • As reported in my elog yesterday, both the left magnets (UL and LL) seemed too low relative to the OSEM coils
  • Eric and I checked the height of the scribe lines using the microscope and found that the scribe lines were low on the left side and high on the right side (as viewed from the AR side) by approximately the same amount, confirming our suspicion that the optic was rotated. The position of the scribe line on the bottom of the optic relative to the bottom-rear face EQ stop also suggested the same
  • Eric brought in the bottom EQ stops, and once the wire was slightly unloaded, rotated the optic by the required amount by hand 
  • This process took two tries, but we were successful
  • Re-checked heights of scribe lines using microscope, and once we were satisfied, re-did the coarse pitch balancing

Part 2: Replacement of holder for top pair of OSEMs

  • Eric and I had difficulty removing the UR OSEM-holding screw
  • This is the non-silver-coated new variety of screw
  • It got to a point where I could neither move the screw in or out, even with the help of a pair of pliers
  • I decided to swap out the piece of the suspension tower holding the top two OSEMs (UR and UL) with the same piece from the old ETMX tower that is currently residing on the flow bench at the south end (along with the accompanying piece that overhangs the optic and holds the front-face and top earthquake stops
  • I cleaned the piece 3-4 times with acetone, and then a couple of times with isopropanol. I adjudged this to be sufficient as we are going to air bake the tower anyways prior to installation in the vacuum chamber
  • I then swapped the pieces:
    • First I brought in the bottom pairs of EQ stops
    • Next, I secured the optic using the three lower face EQ stops
    • Then, I removed the EQ stop screws from the overhanging piece, after which I removed the overhanging piece itself
    • After removing the top-back EQ stop, I removed the OSEM-holding piece from the suspension tower
    • Did the above steps in reverse, installing the new piece
  • All went smoothly. This piece does not have a serial number unfortunately
  • After this, I re-inserted the OSEMs, and judged the magnet-coil alignment to be satisfactory to proceed further
  • We decided to use the old variety of silver plated OSEM holding screws for the top two OSEMs (by choice) and the side OSEM (the new variety is too short anyways). During the course of my work tonight, I found this worked way better. The bottom pair of OSEMs remain held by the new variety of unplated screws. We may want to review whether we really want to use this new type of screws (I believe the idea is to make it easier to tighten and loosen the screws)

Part 3: Fine pitch balancing

  • As per the SOS assembly procedure, I turned off the HEPA filters at the clean bench for this part of the work
  • Checked that the HeNe beam incident on the optic was level with the tabletop, beam height set to 5.5"
  • Proceeded to do the fine pitch balancing the same way as described in yesterday's elog (i.e. no PZT buzzer, just fine touches by hand)
  • I was able to converge fairly quickly to a good point in configuration space
  • After re-centering the OSEM coils such that the PD output was ~50% of its maximum value (see Attachment #1), I found over a lever arm length of 56" (=1.42m) a beam height deviation from 5.5" by <2mm. This corresponds to 0.7mrad pitching forwards towards the HR side
  • The suspension assembly procedure tells us to aim for 0.5mrad, but I think this is close enough for standoff gluing, as this misalignment is extremely sensitive to the OSEM coil positions (although I would say, from Attachment #1, that they are actually pretty well centered)
  • The only thing that concerns me is that the LL magnet is still a little low relative to the coil. This can be fixed by shimming if necessary...

Attachment #1: Striptool trace showing OSEMs are pretty well centered (towards the end, I turned on the HEPA filters again, which explains the shift of the traces). The y-axis is normalized such that the maximum displayed corresponds to the fully open PD output of the coils

Attachment #2: Fine pitch balancing optical lever setup

Attachment #3: Tower assembly

Attachment #4: SIDE OSEM close-up

Attachment #5: UR OSEM close-up

Attachment #6: UL OSEM close-up

Attachment #7: LL OSEM close-up (this is the concerning one)

Attachment #8: LR OSEM close-up


We should also check the following (I forgot and don't want to wear my clean jumpsuit again now to take more photos):

  1. Wire is still in groove
  2. Standoff is sitting on the optic barrel and not on epoxy residue of the guiderod
Attachment 1: OSEMs.PDF  377 kB  | Hide | Hide all
OSEMs.PDF
Attachment 2: IMG_2932.JPG  4.870 MB  | Hide | Hide all
IMG_2932.JPG
Attachment 3: IMG_2934.JPG  6.025 MB  | Hide | Hide all
IMG_2934.JPG
Attachment 4: IMG_2935.JPG  5.284 MB  | Hide | Hide all
IMG_2935.JPG
Attachment 5: IMG_2936.JPG  5.244 MB  | Hide | Hide all
IMG_2936.JPG
Attachment 6: IMG_2937.JPG  5.142 MB  | Hide | Hide all
IMG_2937.JPG
Attachment 7: IMG_2938.JPG  4.847 MB  | Hide | Hide all
IMG_2938.JPG
Attachment 8: IMG_2939.JPG  5.435 MB  | Hide | Hide all
IMG_2939.JPG
ELOG V3.1.3-